# Segre products of cluster algebras

### v1

@@ Abstract
We show that under suitable assumptions the Segre product of two graded cluster algebras has a natural cluster algebra structure.
@@ 1 Introduction
Cluster algebras are a class of combinatorially rich algebras arising in a number of algebraic and geometric contexts (see [FWZ21] and references therein). The additional data of a cluster structure leads to the existence of canonical bases, closely related to the canonical bases arising in Lie theory. Important examples of cluster algebras of this type include coordinate algebras of projective varieties and their various types of cells, e.g. Grassmannians ([Sco06]) and Schubert cells ([GLS11]), positroid varieties and positroid cells ([GL19]), etc.
@@
As we will see, the standard Segre construction imposes significant restrictions on both the graded cluster algebras and the choice of clusters at which one can glue. The latter is perhaps surprising since most cluster algebra constructions are agnostic as to choices of initial cluster.
@@ Acknowledgements
The authors acknowledge financial support from the Engineering and Physical Sciences Research Council (studentship ref. 2436773,https://gtr.ukri.org/projects?ref=studentship-2436773).
@@ 2 Segre Products of Graded Cluster Algebras
It was shown in [GL19] that coordinate rings of positroid varieties in the Grassmannian have cluster algebra structures. In [Pre23], the Segre product of two such cluster algebras is shown to have a cluster structure. In what follows, we aim to generalise this construction to the case of graded skew-symmetric cluster algebras.
@@
  • $\underline{\textup{ex}}_{i}\subsetneq\{1,\dotsc,n_{i}\}$ is the set of indices corresponding to mutable cluster variables;
  • every frozen variable (i.e. those elements with index in $\{1,\dotsc,n_{i}\}\setminus\underline{\textup{ex}}_{i}$) is invertible;
  • $B_{i}$ is an exchange matrix with skew-symmetric principal part;
@@
Let $s_{i}\in\{1,\dotsc,n_{i}\}\setminus\underline{\textup{ex}}_{i}$ be an index corresponding to a frozen cluster variable. We denote by $B_{i}^{s_{i}}$ the $s_{i}$th row of $B_{i}$ and by $\widehat{B}_{i}^{s_{i}}$ the exchange matrix obtained from $B_{i}$ with the $s_{i}$th row removed.
@@
We have also asked that every frozen vertex is invertible, which is a common but not universal assumption in cluster theory. In fact, an examination of our construction shows that this assumption can be weakened to only asking that the glued frozen variables are invertible, which may be a more appropriate assumption for geometric applications.
@@ 2.1 A gluing construction
Fix $s_{1}\in\{1,\dotsc,n_{1}\}\setminus\underline{\textup{ex}}_{1}$ and $s_{2}\in\{1,\dotsc,n_{2}\}\setminus\underline{\textup{ex}}_{2}$ such that $(G_{1})_{s_{1}}=(G_{2})_{s_{2}}$. We will identify the frozen variables $x_{s_{1}}$ and $y_{s_{2}}$, denoting a new proxy variable replacing both of these by $x_{s}$.
@@
$\underline{x}=(\underline{x}_{1}\setminus\{x_{s_{1}}\})\cup(\underline{x}_{2}\setminus\{y_{s_{2}}\})\cup\{x_{s}\}.$
The index set corresponding to mutable variables is given by $\underline{\textup{ex}}=\underline{\textup{ex}}_{1}\cup\underline{\textup{ex}}_{2}$, and for the initial exchange matrix, we take
$\underline{B}=\begin{bmatrix}\widehat{B}_{1}^{s_{1}}&0\\ 0&\widehat{B}_{2}^{s_{2}}\\ B_{1}^{s_{1}}&B_{2}^{s_{2}}\end{bmatrix}.$
@@
$\underline{G}=((\widehat{G}_{1}^{s_{1}})^{T}\ (\widehat{G}_{2}^{s_{2}})^{T}\ G_{1}^{s})^{T}$
where $\widehat{G}_{i}^{s_{i}}$ is the grading vector $G_{i}$ with the $s_{i}$th entry removed and $G_{1}^{s}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}G^{1}_{s_{1}}=G^{2}_{s_{2}}$. We can now define a cluster algebra $\mathcal{A}_{1}\Box\mathcal{A}_{2}=\mathcal{A}(\underline{x},\underline{\textup{ex}},\underline{B},\underline{G})$ from this initial data.
@@
Let $\mathcal{A}_{1}=(\underline{x}_{1}=\{x_{1},x_{2},x_{3}\},\underline{\textup{ex}}_{1}=\{1\},Q_{1},G_{1}=\mathds{1})$ and $\mathcal{A}_{2}=(\underline{x}_{2}=\{y_{1},y_{2},y_{3}\},\underline{\textup{ex}}_{2}=\{1\},Q_{2},G_{1}=\mathds{1})$ be cluster algebras with exchange quivers as follows:
@@
The quiver obtained by ‘gluing’ at the frozen variables $x_{3}$ and $y_{3}$ is shown below—we denote the new variable by $x_{s}$.
@@
$(\underline{x}=\{x_{1},x_{2},y_{1},y_{2},x_{s}\},\underline{\textup{ex}}=\{1,3\},Q,G=\mathds{1}).$
@@
Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be graded cluster algebras. Fix $s_{1}\in\{1,\dotsc,n_{1}\}\setminus\underline{\textup{ex}}_{1}$ and $s_{2}\in\{1,\dotsc,n_{2}\}\setminus\underline{\textup{ex}}_{2}$ such that $(G_{1})_{s_{1}}=(G_{2})_{s_{2}}$. Then the cluster algebras $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ and $\mathcal{A}_{2}\Box\mathcal{A}_{1}$ are isomorphic as cluster algebras.
@@
Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be graded cluster algebras. Fix $s_{1}\in\{1,\dotsc,n_{1}\}\setminus\underline{\textup{ex}}_{1}$ and $s_{2}\in\{1,\dotsc,n_{2}\}\setminus\underline{\textup{ex}}_{2}$ such that $(G_{1})_{s_{1}}=(G_{2})_{s_{2}}$.
@@
This follows from observing that our gluing process does not introduce any new arrows between mutable vertices. Since mutation is a local phenomenon and concentrated on mutable vertices, it is straightforward to see that mutating at vertices indexed by $\underline{\textup{ex}}_{1}$ is independent of mutating at vertices indexed by $\underline{\textup{ex}}_{2}$ and the variables obtained are exactly as if the gluing had not been carried out.
There is an overall reduction of one in the number of cluster variables because we have glued two previously distinct frozen variables; note that this highlights the difference between this construction and the usual “disconnected” product of cluster algebras (where one simply takes the union of clusters and direct sum of exchange matrices). ∎
@@ 2.2 Relationship with the Segre product
The cluster algebra $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ defined above does not have any immediately obvious relationship with the Segre product $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$. Our main goal in what follows is to show that there indeed is one and furthermore, we will establish under what conditions these algebras are actually isomorphic.
We first identify a candidate isomorphism, analogous to the map $\delta^{\text{src}}$ defined in [Pre23].
Proposition 2.6.
Let $\mathcal{A}_{i}=(\underline{x}_{i},\underline{\textup{ex}}_{i},B_{i},G_{i})$, $i=1,2$ be graded cluster algebras and fix $s_{1}\in\{1,\dotsc,n_{1}\}\setminus\underline{\textup{ex}}_{1}$ and $s_{2}\in\{1,\dotsc,n_{2}\}\setminus\underline{\textup{ex}}_{2}$.
Then the map $\varphi:\mathcal{A}_{1}\Box\mathcal{A}_{2}\to\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ given on initial cluster variables by
$\varphi(x_{j})$ $=x_{j}\otimes y_{s_{2}}^{\deg x_{j}}\quad\text{for }j\in\{1,\dotsc,n_{1}\},$
$\varphi(y_{j})$ $=x_{s_{1}}^{\deg y_{j}}\otimes y_{j}\quad\text{for }j\in\{1,\dotsc,n_{2}\}\ \text{and}$
$\varphi(x_{s})$ $=x_{s_{1}}\otimes y_{s_{2}}$
is an injective algebra homomorphism, with the property that the above formulæ hold for any cluster of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$.
@@
Let $\varphi$ denote the algebra homomorphism between fields of rational functions
$\varphi\colon\mathbb{K}(\underline{x})\to\mathbb{K}(\underline{x}_{1}\otimes y_{s_{2}}^{\deg x_{\bullet}},x_{s_{1}}^{\deg y_{\bullet}}\otimes\underline{x}_{2},x_{s_{1}}\otimes y_{s_{2}})$
obtained from the above specification on generators of the domain, where $\underline{x}_{1}\otimes y_{s_{1}}^{\deg x_{\bullet}}=\{x\otimes y_{s_{2}}^{\deg x}\mid x\in\underline{x}_{1}\}$ and similarly for $x_{s_{1}}\otimes\underline{x}_{2}$.
We claim that the map $\varphi$ has a natural extension to all cluster variables via cluster mutation. To prove this, we proceed by induction on the number of mutation steps from the initial cluster. We compute $\varphi$ for a one-step mutation from the initial cluster as follows.
We first consider the case in which $x_{k}\in\underline{x}_{1}$, $k\in\underline{\textup{ex}}_{1}$. To reduce the proliferation of subscripts, we will write $B^{i}_{jk}$ for the $(j,k)$-entry of $B_{i}$ and $G^{i}_{j}$ for the $j$th entry of $G_{i}$. Note that $G^{1}_{j}=\deg x_{j}$ and $G^{2}_{j}=\deg y_{j}$. We also set $[n]_{+}=\max\{n,0\}$ and $[n]_{-}=\max\{-n,0\}$.
@@
$\varphi(\mu_{k}(x_{k}))$ $=\varphi\Bigg{(}\frac{1}{x_{k}}\Bigg{[}x_{s_{1}}^{[B_{n_{1}+n_{2}-1,k}]_{+}}\Bigg{(}\prod\limits_{B_{jk}>0}x_{j}^{B_{jk}}\Bigg{)}\Bigg{(}\prod\limits_{B_{n_{1}+j-1,k}>0}y_{j}^{B_{n_{1}+j-1,k}}\Bigg{)}$
$\phantom{=}+x_{s_{1}}^{[B_{n_{1}+n_{2}-1,k}]_{-}}\Bigg{(}\prod\limits_{B_{jk}<0}x_{j}^{-B_{jk}}\Bigg{)}\Bigg{(}\prod\limits_{B_{n_{1}+j-1,k}<0}y_{j}^{-B_{n_{1}+j-1,k}}\Bigg{)}\Bigg{]}\Bigg{)}$
$=\varphi\Bigg{(}\frac{1}{x_{k}}\Bigg{[}x_{s_{1}}^{[B_{n_{1}+n_{2}-1,k}]_{+}}\Bigg{(}\prod\limits_{B_{jk}>0}x_{j}^{B_{jk}}\Bigg{)}$
$\phantom{=}+x_{s_{1}}^{[B_{n_{1}+n_{2}-1,k}]_{-}}\Bigg{(}\prod\limits_{B_{jk}<0}x_{j}^{-B_{jk}}\Bigg{)}\Bigg{]}\Bigg{)}$
$=\varphi\Bigg{(}\frac{1}{x_{k}}\Bigg{[}\prod\limits_{B^{1}_{jk}>0}x_{j}^{B^{1}_{jk}}+\prod\limits_{B^{1}_{jk}<0}x_{j}^{-B^{1}_{jk}}\Bigg{]}\Bigg{)}$
$=\frac{1}{x_{k}\otimes y_{s_{2}}^{G_{k}^{1}}}\Bigg{[}\prod\limits_{B^{1}_{jk}>0}\Bigg{(}x_{j}^{B^{1}_{jk}}\otimes y_{s_{2}}^{G_{j}^{1}B^{1}_{jk}}\Bigg{)}+\prod\limits_{B^{1}_{jk}<0}\Bigg{(}x_{j}^{-B^{1}_{jk}}\otimes y_{s_{2}}^{-G_{j}^{1}B^{1}_{jk}}\Bigg{)}\Bigg{]}$
$=\frac{1}{x_{k}\otimes y_{s_{2}}^{G_{k}^{1}}}\Bigg{[}\prod\limits_{B^{1}_{jk}>0}\Bigg{(}x_{j}^{B^{1}_{jk}}\Bigg{)}\otimes y_{s_{2}}^{d}+\prod\limits_{B^{1}_{jk}<0}\Bigg{(}x_{j}^{-B^{1}_{jk}}\Bigg{)}\otimes y_{s_{2}}^{d}\Bigg{]}$
$=\frac{1}{x_{k}}\Bigg{(}\prod\limits_{B^{1}_{jk}>0}x_{j}^{B^{1}_{jk}}+\prod\limits_{B^{1}_{jk}<0}x_{j}^{-B^{1}_{jk}}\Bigg{)}\otimes y_{s_{2}}^{d-\deg x_{k}}$
$=\mu_{k}(x_{k})\otimes y_{s_{2}}^{d-\deg x_{k}}$
$=\mu_{k}(x_{k})\otimes y_{s_{2}}^{\deg\mu_{k}(x_{k})}$

### v2

@@ Abstract
We show that under mild assumptions the Segre product of two graded cluster algebras has a natural cluster algebra structure.
MSC:
13F60
@@ 1 Introduction
Cluster algebras are a class of combinatorially rich algebras arising in a number of algebraic and geometric contexts (see [FWZ21] and references therein). The additional data of a cluster structure leads to the existence of canonical bases, closely related to the canonical bases arising in Lie theory. Important examples of cluster algebras of this type include coordinate algebras of projective varieties and their various types of cells, e.g. Grassmannians ([Sco06]), Schubert cells ([GLS11]) and positroid varieties ([GL19]).
@@ Acknowledgements
The authors acknowledge financial support from the Engineering and Physical Sciences Research Council (studentship ref. 2436773,https://gtr.ukri.org/projects?ref=studentship-2436773). We are also very grateful to the anonymous referee for many helpful suggestions that have greatly improved the exposition of our results.
@@ 2 Segre Products of Graded Cluster Algebras
It was shown by Galashin and Lam in [GL19] that coordinate rings of positroid varieties in the Grassmannian have cluster algebra structures. This class is closed under Segre product and in [Pre23], Pressland shows how the Galashin–Lam cluster structure on the product is related to that on the factors.
In what follows, we aim to generalise this construction to the case of graded skew-symmetric cluster algebras: we start with two graded cluster algebras and show that their Segre product has a natural cluster structure. For coordinate rings of positroid varieties, Pressland’s result shows that the Galashin–Lam cluster structure on the product is equal to that obtained by the Segre product construction we give here.
@@
  • $\underline{\textup{ex}}_{i}\subsetneq\underline{x}_{i}$ is the set of mutable cluster variables;
  • every frozen variable (i.e. those elements in $\underline{x}_{i}\setminus\underline{\textup{ex}}_{i}$) is invertible;
  • $B_{i}$ is an exchange matrix (with rows indexed by $\underline{x}_{i}$ and columns by $\underline{\textup{ex}}_{i}$) with skew-symmetric principal part;
@@
It is common to visualise cluster mutation using quivers; see e.g. [Mar14]. We will do the same: an exchange matrix $B_{i}$ will be represented by an ice quiver having vertices $\bullet$ labelled by the elements of $\underline{x}_{i}$. The frozen variables indicated by a box $\Box$, to indicate that mutation is not carried out there. Arrows are determined by $B_{i}$: the number of arrows from $x_{j}$ to $x_{k}$ is $(B_{i})_{jk}$.
@@
Let $\underline{x}$ be a cluster with $x$ a cluster variable and $B$ the exchange matrix associated to $\underline{x}$. We denote by $B^{x}$ the row of $B$ indexed by $x$ and by $\widehat{B}^{x}$ the matrix obtained from $B$ by removing the row $B^{x}$. If $x$ is frozen, $\widehat{B}^{x}$ is again an exchange matrix.
@@
We have also asked that every frozen variable is invertible, which is a common but not universal assumption in cluster theory. In fact, an examination of our construction shows that this assumption can be weakened to only asking that the glued frozen variables are invertible, which may be a more appropriate assumption for geometric applications.
@@ 2.1 A gluing construction
Fix $x\in\underline{x}_{1}\setminus\underline{\textup{ex}}_{1}$ and $y\in\underline{x}_{2}\setminus\underline{\textup{ex}}_{2}$ such that $(G_{1})_{x}=(G_{2})_{y}$. That is, $x$ and $y$ are frozen variables in their respective clusters and their degrees are equal. We will identify the frozen variables $x$ and $y$, denoting a new proxy variable replacing both of these by $z$.
@@
$\underline{x}_{1}\Box\underline{x}_{2}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}(\underline{x}_{1}\setminus\{x\})\cup(\underline{x}_{2}\setminus\{y\})\cup\{z\}.$
The mutable variables are $\underline{\textup{ex}}_{1}\cup\underline{\textup{ex}}_{2}$, and for the initial exchange matrix, we form the block matrix
$B_{1}\Box B_{2}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{bmatrix}\widehat{B}_{1}^{x}&0\\ 0&\widehat{B}_{2}^{y}\\ B_{1}^{x}&B_{2}^{y}\end{bmatrix}.$
@@
$G_{1}\Box G_{2}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{bmatrix}\widehat{G}_{1}^{x}\\ \widehat{G}_{2}^{y}\\ G_{1}^{z}\end{bmatrix}$
where $\widehat{G}_{1}^{x}$ is the grading vector $G_{1}$ with the entry indexed by $x$ removed (and similarly for $\widehat{G}_{2}^{y}$) and $G_{1}^{z}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}(G_{1})_{x}=(G_{2})_{y}$. We can now define a cluster algebra
$\mathcal{A}_{1}\Box\mathcal{A}_{2}=\mathcal{A}(\underline{x}_{1}\Box\underline{x}_{2},\underline{\textup{ex}}_{1}\cup\underline{\textup{ex}}_{2},B_{1}\Box B_{2},G_{1}\Box G_{2})$
from this initial data.
Let us extend the above notation to write
$\underline{x}_{1}^{\prime}\Box\underline{x}_{2}^{\prime}=(\underline{x}_{1}^{\prime}\setminus\{x\})\cup(\underline{x}_{2}^{\prime}\setminus\{y\})\cup\{z\},$
where $\underline{x}_{1}^{\prime}$, $\underline{x}_{2}^{\prime}$ are now allowed to be any clusters of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively, and say that $\underline{x}_{1}^{\prime}\Box\underline{x}_{2}^{\prime}$ is obtained by gluing $x$ and $y$. This is well-defined since $x$ and $y$ are frozen. Similarly, we extend the notation $B_{1}\Box B_{2}$ and $G_{1}\Box G_{2}$ to any appropriate input matrices/vectors.
@@
Let $\mathcal{A}_{1}=(\underline{x}_{1}=\{x_{1},x_{2},x_{3}\},\underline{\textup{ex}}_{1}=\{x_{1}\},Q_{1},G_{1}=\mathds{1})$ and $\mathcal{A}_{2}=(\underline{x}_{2}=\{y_{1},y_{2},y_{3}\},\underline{\textup{ex}}_{2}=\{y_{1}\},Q_{2},G_{1}=\mathds{1})$ be cluster algebras with exchange quivers as follows:
@@
The quiver obtained by ‘gluing’ at the frozen variables $x_{3}$ and $y_{3}$ is shown below—we denote the new variable by $z$.
@@
$(\underline{x}=\{x_{1},x_{2},y_{1},y_{2},z\},\underline{\textup{ex}}=\{x_{1},y_{1}\},Q,G=\mathds{1}).$
@@
Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be graded cluster algebras. Fix $x\in\underline{x}_{1}\setminus\underline{\textup{ex}}_{1}$ and $y\in\underline{x}_{2}\setminus\underline{\textup{ex}}_{2}$ such that $(G_{1})_{x}=(G_{2})_{y}$. Then the cluster algebras $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ and $\mathcal{A}_{2}\Box\mathcal{A}_{1}$ are isomorphic as cluster algebras.
@@
Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be graded cluster algebras. Fix $x\in\underline{x}_{1}\setminus\underline{\textup{ex}}_{1}$ and $y\in\underline{x}_{2}\setminus\underline{\textup{ex}}_{2}$ such that $(G_{1})_{x}=(G_{2})_{y}$.
  1. (i)
    Every cluster variable of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ is naturally identified with a cluster variable of $\mathcal{A}_{1}$, a cluster variable of $\mathcal{A}_{2}$ or is equal to $z$.
  2. (ii)
    There is a bijection between pairs of clusters $(\underline{x}_{1}^{\prime},\underline{x}_{2}^{\prime})$ and clusters of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ given by gluing, i.e. sending $(\underline{x}_{1}^{\prime},\underline{x}_{2}^{\prime})$ to $\underline{x}_{1}^{\prime}\Box\underline{x}_{2}^{\prime}$ for a cluster $\underline{x}_{1}^{\prime}$ of $\mathcal{A}_{1}$ and $\underline{x}_{2}^{\prime}$ of $\mathcal{A}_{2}$.
Proof.
This follows from observing that our gluing process does not introduce any new arrows between mutable vertices. Since mutation is a local phenomenon and concentrated on mutable vertices, it is straightforward to see that mutating at vertices indexed by $\underline{\textup{ex}}_{1}$ is independent of mutating at vertices indexed by $\underline{\textup{ex}}_{2}$ and the (mutable) variables obtained are exactly as if the gluing had not been carried out. The frozen variables of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ are those of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ excluding $x$ and $y$, along with the glued frozen $z$.
For the second part, note that the same argument shows that there is a similar bijection for the clusters of $\mathcal{A}_{1}\times\mathcal{A}_{2}$, where the latter denotes the “disconnected” product of cluster algebras, where one simply takes the union of clusters and direct sum of exchange matrices. Now there is evidently a bijection between the clusters of $\mathcal{A}_{1}\times\mathcal{A}_{2}$ and those of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$, given by $\underline{x}_{1}^{\prime}\cup\underline{x}_{2}^{\prime}\mapsto\underline{x}_{1}^{\prime}\Box\underline{x}_{2}^{\prime}$, from which the claim follows. ∎
Corollary 2.5.
Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be graded cluster algebras. Fix $x\in\underline{x}_{1}\setminus\underline{\textup{ex}}_{1}$ and $y\in\underline{x}_{2}\setminus\underline{\textup{ex}}_{2}$ such that $(G_{1})_{x}=(G_{2})_{y}$.
@@
These are now immediate from the previous lemma. Note that there is an overall reduction of one in the number of cluster variables because we have glued two previously distinct frozen variables; this highlights the difference between this construction and the disconnected product. ∎
@@ 2.2 Relationship with the Segre product
Our main result is the following theorem, showing that the cluster algebra construction $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ induces a cluster algebra structure on the Segre product. The isomorphism we will use is directly analogous to the map $\delta^{\text{src}}$ defined in [Pre23].
Theorem 2.7.
Let $\mathcal{A}_{i}=(\underline{x}_{i},\underline{\textup{ex}}_{i},B_{i},G_{i})$, $i=1,2$ be graded cluster algebras such that there exist $x\in\underline{x}_{1}\setminus\underline{\textup{ex}}_{1}$ and $y\in\underline{x}_{2}\setminus\underline{\textup{ex}}_{2}$ both of degree 1.
$\varphi(x_{j})$ $=x_{j}\otimes y^{\deg x_{j}}\quad\text{for }x_{j}\in\underline{x}_{1}\setminus\{x\},$
$\varphi(y_{j})$ $=x^{\deg y_{j}}\otimes y_{j}\quad\text{for }y_{j}\in\underline{x}_{2}\setminus\{y\}\ \text{and}$
$\varphi(z)$ $=x\otimes y$
@@
Recalling that we set
$\underline{x}_{1}\Box\underline{x}_{2}=(\underline{x}_{1}\setminus\{x\})\cup(\underline{x}_{2}\setminus\{y\})\cup\{z\},$
let $\varphi$ denote the algebra homomorphism $\varphi\colon\mathbb{K}(\underline{x}_{1}\Box\underline{x}_{2})\to\mathbb{K}(\underline{x}_{1})\otimes\mathbb{K}(\underline{x}_{2})$ obtained from the above specification on generators of the domain. This map is injective since the elements $\varphi(x_{j})$, $\varphi(y_{j})$ and $\varphi(z)$ are algebraically independent.
Now let $\varphi$ denote the restriction of the above map to $\mathcal{A}_{1}\Box\mathcal{A}_{2}$. We first claim that the restricted map $\varphi$ takes values in the subalgebra $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$. To prove this, we proceed by induction on the number of mutation steps from the initial cluster.
We may take as base case that of zero mutations from the initial cluster: there is nothing to do, since we see immediately that $\varphi(x_{j})$, $\varphi(y_{j})$ and $\varphi(z)$ lie in $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ by definition.
Now assume that the result holds $r-1$ mutations from the initial cluster $\underline{x}=\underline{x}_{1}\Box\underline{x}_{2}$ (for $r\geq 1$) of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$. That is, let $\underline{y}=\mu_{k_{r-1}}\mu_{k_{r-2}}\dotsm\mu_{k_{1}}(\underline{x})$. Set $B=\mu_{k_{r-1}}\mu_{k_{r-2}}\dotsm\mu_{k_{1}}(B_{1}\Box B_{2})$.
By Lemma 2.4(ii), we have that $\underline{y}=\underline{y}_{1}\Box\underline{y}_{2}$ for some clusters $\underline{y}_{1}$, $\underline{y}_{2}$ of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively. Moreover, there is a decompostion
$\{k_{1},\dotsc,k_{r-1}\}=\{l_{1},\dotsc,l_{s}\}\sqcup\{m_{1},\dotsc,m_{t}\}$
such that $\underline{y}_{1}=\mu_{l_{s}}\dotsm\mu_{l_{1}}(\underline{x}_{1})$ and $\underline{y}_{2}=\mu_{m_{t}}\dotsm\mu_{m_{1}}(\underline{x}_{2})$.
Let $\underline{y}_{1}=\{x_{1},\dotsc,x_{n_{1}}\}$ and $\underline{y}_{2}=\{y_{1},\dotsc,y_{n_{2}}\}$, so that
$\underline{y}=\underline{y}_{1}\Box\underline{y}_{2}=(\{x_{1},\dotsc,x_{n_{1}}\}\setminus\{x\})\sqcup(\{y_{1},\dotsc,y_{n_{2}}\}\setminus\{y\})\sqcup\{z\}$
Let $C=\mu_{l_{s}}\dotsm\mu_{l_{1}}(B_{1})$, $D=\mu_{m_{t}}\dotsm\mu_{m_{1}}(B_{2})$, $H=\mu_{l_{s}}\dotsm\mu_{l_{1}}(G_{1})$ and $K=\mu_{m_{t}}\dotsm\mu_{m_{1}}(G_{2})$. Then in particular $B=C\Box D$ and $H_{j}=\deg x_{j}$ and $K_{j}=\deg y_{j}$. We also set $[n]_{+}=\max\{n,0\}$ and $[n]_{-}=\max\{-n,0\}$.
We then compute $\varphi$ for one further mutation in direction $k_{r}=k$. We first consider the case in which $x_{k}\in\underline{y}_{1}$ is mutable.
@@
$\varphi(\mu_{k}(x_{k}))$ $=\varphi\Bigg{(}\frac{1}{x_{k}}\Bigg{[}\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}x_{j}^{[B_{x_{j},x_{k}}]_{+}}\Bigg{)}\Bigg{(}\prod\limits_{y_{j}\in\underline{y}_{2}\setminus\{y\}}y_{j}^{[B_{y_{j},x_{k}}]_{+}}\Bigg{)}z^{[B_{z,x_{k}}]_{+}}$
$\qquad\qquad+\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}x_{j}^{[B_{x_{j},x_{k}}]_{-}}\Bigg{)}\Bigg{(}\prod\limits_{y_{j}\in\underline{y}_{2}\setminus\{y\}}y_{j}^{[B_{y_{j},x_{k}}]_{-}}\Bigg{)}z^{[B_{z,x_{k}}]_{-}}\Bigg{]}\Bigg{)}$
$=\varphi\Bigg{(}\frac{1}{x_{k}}\Bigg{[}\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}x_{j}^{[B_{x_{j},x_{k}}]_{+}}\Bigg{)}z^{[B_{z,x_{k}}]_{+}}+\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}x_{j}^{[B_{x_{j},x_{k}}]_{-}}\Bigg{)}z^{[B_{z,x_{k}}]_{-}}\Bigg{]}\Bigg{)}$
$=\varphi\Bigg{(}\frac{1}{x_{k}}\Bigg{[}\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}x_{j}^{[C_{x_{j},x_{k}}]_{+}}\Bigg{)}z^{[C_{x,x_{k}}]_{+}}+\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}x_{j}^{[C_{x_{j},x_{k}}]_{-}}\Bigg{)}z^{[C_{x,x_{k}}]_{-}}\Bigg{]}\Bigg{)}$
$=\frac{1}{x_{k}\otimes y^{\deg x_{k}}}\Bigg{[}\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}\Bigg{(}x_{j}^{[C_{x_{j},x_{k}}]_{+}}\otimes y^{[C_{x_{j},x_{k}}]_{+}\deg x_{j}}\Bigg{)}x^{[C_{x,x_{k}}]_{+}}\otimes y^{[C_{x,x_{k}}]_{+}}$
$\phantom{=}+\prod\limits_{x_{j}\in\underline{y}_{1}\setminus\{x\}}\Bigg{(}x_{j}^{[C_{x_{j},x_{k}}]_{-}}\otimes y^{[C_{x_{j},x_{k}}]_{-}\deg x_{j}}\Bigg{)}x^{[C_{x,x_{k}}]_{-}}\otimes y^{[C_{x,x_{k}}]_{-}}\Bigg{]}$
$=\frac{1}{x_{k}\otimes y^{\deg x_{k}}}\Bigg{[}\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{+}}\otimes y^{d}+\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{-}}\otimes y^{d}\Bigg{]}$
$=\frac{1}{x_{k}}\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{+}}+\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{-}}\Bigg{)}\otimes y^{d-\deg x_{k}}$
$=\mu_{k}(x_{k})\otimes y^{d-\deg x_{k}}$
$=\mu_{k}(x_{k})\otimes y^{\deg\mu_{k}(x_{k})}$